OpenStudy (anonymous):

solve for x: |x+2|-10= 5

OpenStudy (anonymous):

Add 10 to both sides. |x+2| = 15 Now you have two equations x+2 = 15 and x+2 = -15 Solve for x, and you get x = 13 or x = -17

OpenStudy (katrinakaif):

|x+2| = 5 + 10 |x+2| = 15 x + 2 = 15 x +2 = -15 and solve for both x values..I believe

OpenStudy (anonymous):

agh, i meant |x+2|+10=5, not minus 10.

OpenStudy (anonymous):

Subtract 10 from both sides, |x+2|=-5 Split into two equations: x+2=-5 x+2=5 x= -7 or x = 3